Betty bought a car which was $1543.00 more than the other car that Joan bought. Paula bought a car that cost $977.00 less than Joans. The total price of all three cars was $62468.00 How much did each girl's car cost?
Answer by oberobic(2304)

In the setup it is important to use the fewest possible unknowns.
B is the price paid by Betty for her car.
P is the price paid by Paula for her car.
J is the price paid by Joan for her car.
B + P + J = 62468.
B = J + 1543 (Betty's car cost 1543 more than Joan's)
P = J - 977 (Paula's car is 977 less than Joan's)
Substituting in the sum formula, we have:
(J+1543) + (J-977) + J = 62468
Combining...
3J + 566 = 62468
Subtracting 566 from both sides...
3J = 61,902
Dividing both sides by 3...
J = 20,634
B = 20,634 + 1543
P = 20,634 - 977